OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (amistre64):

write f(g), then replace g with its definition

OpenStudy (amistre64):

then simplify to your hearts desire

OpenStudy (anonymous):

i got 74x-6

OpenStudy (anonymous):

-40x + 24 -40x + 31 -40x + 64 -40x + 70 are the answers

OpenStudy (anonymous):

\[f(g(x))=f(10x-6)=-4(10x-6)+7\] is a start

OpenStudy (anonymous):

multiply and combine like terms to finish

OpenStudy (anonymous):

@amistre64 Hello!!

OpenStudy (amistre64):

f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). f(g) = -4g + 7, and g is defined already f(g(x)) = -4(10x-6) + 7

OpenStudy (amistre64):

howdy :)

OpenStudy (anonymous):

got it, thank you